Literatur

[1] Alexander, J. W. : Topological invariants of knots and links. In: Trans. Amer. Math. Soc. 30 (1928), Nr. 2, S. 275–306. – ISSN 0002–9947

[2] Báyer, P. ; Neukirch, J. : On values of zeta functions and l-adic Euler characteristics. In: Invent. Math. 50 (1978/79), Nr. 1, S. 35–64. – ISSN 0020–9910

[3] Bolle, L. : Knoten. 6. Auflage. Edition Maritim, 2002. – Mit über 100 praxisgerechten Fotos. – ISBN 3–89225–354–4

[4] Bourbaki, N. : Commutative algebra. Chapters 1–7. Berlin: Springer-Verlag, 1998 (Elements of Mathematics). – Translated from the French, Reprint of the 1989 English translation. – ISBN 3–540–64239–0

[5] Brown, K. S. : Cohomology of groups. New York: Springer-Verlag, 1982. – Graduate Texts in Mathematics, Bd. 87. – ISBN 0–387–90688–6

[6] Burde, G. ; Zieschang, H. : Knots. Berlin: Walter de Gruyter & Co., 1985 (de Gruyter Studies in Mathematics, Bd. 5). – ISBN 3–11–008675–1

[7] Crowell, R. H. : Corresponding group and module sequences. In: Nagoya Math. J. 19 (1961), S. 27–40

[8] Deninger, C. : A note on arithmetic topology and dynamical systems. In: Algebraic number theory and algebraic geometry Bd. 300. Providence, RI: Amer. Math. Soc., 2002, S. 99–114

[9] Eisenbud, D. : Commutative algebra. With a view toward algebraic geometry. New York: Springer-Verlag, 1995. – Graduate Texts in Mathematics, Bd. 150. – ISBN 0–387– 94268–8; 0–387–94269–6

[10] Fox, R. H. : Free differential calculus. I. Derivation in the free group ring. In: Ann. of Math. (2) 57 (1953), S. 547–560

[11] Fox, R. H. : Free differential calculus. II. The isomorphism problem of groups. In: Ann. of Math. (2) 59 (1954), S. 196–210

[12] Gordon, C. M. ; Luecke, J. : Knots are determined by their complements. In: J. Amer. Math. Soc. 2 (1989), Nr. 2, S. 371–415. – ISSN 0894–0347

[13] Grothendieck, A. : Revêtements étales et groupe fondamental. Berlin: Springer-Verlag, 1971. – Séminaire de Géométrie Algébrique du Bois Marie 1960–1961 (SGA 1), Lecture Notes in Mathematics, Vol. 224

[14] Hartshorne, R. : Algebraic geometry. New York: Springer-Verlag, 1977. – Graduate Texts in Mathematics, Bd. 52. – ISBN 0–387–90244–9

[15] Hatcher, A. : Algebraic topology. Cambridge: Cambridge University Press, 2002. – ISBN 0–521–79160–X; 0–521–79540–0

[16] Iwasawa, K. : On Γ-extensions of algebraic number fields. In: Bull. Amer. Math. Soc. 65 (1959), S. 183–226 [17] Iwasawa, K. : Lectures on p-adic L-functions. Princeton, New Jersey: Princeton Uni-

versity Press, 1972. – ISBN 0–691–08112–3

[18] Iwasawa, K. : On ℤ_{l}-extensions of algebraic number fields. In: Ann. of Math. (2) 98 (1973), S. 246–326

[19] Kawauchi, A. : A survey of knot theory. Basel: Birkhäuser Verlag, 1996. – Translated and revised from the 1990 Japanese original by the author. – ISBN 3–7643–5124–1

[20] Koblitz, N. : p-adic numbers, p-adic analysis, and zeta-functions. Second edition. New York: Springer-Verlag, 1984. – Graduate Texts in Mathematics, Bd. 58. – ISBN 0–387– 96017–1

[21] Koch, H. : Galoissche Theorie der p-Erweiterungen. Berlin: Springer-Verlag, 1970. – Mit einem Geleitwort von I. R. Šafarevič

[22] Koch, H. : l-Erweiterungen mit vorgegebenen Verzweigungsstellen. In: J. Reine Angew. Math. 219 (1965), S. 30–61

[23] Lang, S. : Algebra. 3rd edition. New York: Springer-Verlag, 2002. – Graduate Texts in Mathematics, Bd. 211. – ISBN 0–387–95385–X

[24] Lickorish, W. B. R. : An introduction to knot theory. New York: Springer-Verlag, 1997. – Graduate Texts in Mathematics, Bd. 175. – ISBN 0–387–98254–X

[25] Long, R. L. : Algebraic number theory. New York: Marcel Dekker Inc., 1977. – Monographs and Textbooks in Pure and Applied Mathematics, Vol. 41. – ISBN 0–8247–6540–0

[26] Lyndon, R. C. : Cohomology theory of groups with a single defining relation. In: Ann. of Math. (2) 52 (1950), S. 650–665

[27] Mazur, B. : Notes on étale cohomology of number fields. In: Ann. Sci. École Norm. Sup. (4) 6 (1973), S. 521–552 (1974)

[28] Milne, J. S. : Arithmetic duality theorems. Boston, MA: Academic Press Inc., 1986. – Perspectives in Mathematics, Bd. 1. – ISBN 0–12–498040–6

[29] Milne, J. S. : Étale cohomology. Princeton, N.J.: Princeton University Press, 1980. – Princeton Mathematical Series, Bd. 33. – ISBN 0–691–08238–3

[30] Murre, J. P. : Lectures on an introduction to Grothendieck’s theory of the fundamental group. Bombay: Tata Institute of Fundamental Research, 1967. – Notes by S. Anantharaman, Tata Institute of Fundamental Research Lectures on Mathmatics, No 40

[31] Neukirch, J. : Algebraische Zahlentheorie. In: Ein Jahrhundert Mathematik 1890–1990 Bd. 6. Braunschweig: Vieweg, 1990, S. 587–628

[32] Neukirch, J. : Algebraic number theory. Berlin: Springer-Verlag, 1999 (Grundlehren der Mathematischen Wissenschaften, Bd. 322). – Translated from the 1992 German original.

– ISBN 3–540–65399–6 [33] Neukirch, J. ; Schmidt, A. ; Wingberg, K. : Cohomology of number fields. Berlin:

Springer-Verlag, 2000 (Grundlehren der Mathematischen Wissenschaften, Bd. 323). – ISBN 3–540–66671–0

[34] Nguy~ ^en-Quang-— D~ ^o, T. : Formations de classes et modules d’Iwasawa. In: Number theory, Noordwijkerhout 1983. Berlin: Springer, 1984. – Lecture Notes in Mathematics, Bd. 1068, S. 167–185

[35] Ramachandran, N. : A note on arithmetic topology. In: C. R. Math. Acad. Sci. Soc. R. Can. 23 (2001), Nr. 4, S. 130–135. – ISSN 0706–1994

[36] Raynaud, M. : Anneaux locaux henséliens. Berlin: Springer-Verlag, 1970 (Lecture Notes in Mathematics, Vol. 169)

[37] Reznikov, A. : Hakenness and b_{1}. – Preprint 1998, www.arxiv.org/abs/math.DG/9903064

[38] Reznikov, A. : Embedded incompressible surfaces and homology of ramified coverings of three-manifolds. In: Selecta Math. (N.S.) 6 (2000), Nr. 1, S. 1–39. – ISSN 1022–1824

[39] Ribes, L. ; Zalesskii, P. : Profinite groups. Berlin: Springer-Verlag, 2000 (Ergebnisse der Mathematik und ihrer Grenzgebiete. 3. Folge. A Series of Modern Surveys in Mathematics, Bd. 40). – ISBN 3–540–66986–8

[40] Serre, J.-P. : Cohomologie galoisienne. Cinquième édition. Berlin: Springer-Verlag, 1994 (Lecture Notes in Mathematics, Bd. 5). – ISBN 3–540–58002–6

[41] Serre, J.-P. : Classes des corps cyclotomiques (d’après K. Iwasawa). In: Séminaire Bourbaki, Vol. 5,. Paris: Soc. Math. France, 1995, Exp. No. 174, S. 83–93

[42] Sikora, A. S. : Analogies between group actions on 3-manifolds and number fields. – Preprint 2003 (revised v2), www.arxiv.org/abs/math.GT/0107210

[43] Stark, H. M. : A complete determination of the complex quadratic fields of class-number one. In: Michigan Math. J. 14 (1967), S. 1–27

[44] Trotter, H. F. : Homology of group systems with applications to knot theory. In: Ann. of Math. (2) 76 (1962), S. 464–498

[45] Turaev, V. : Introduction to combinatorial torsions. Basel: Birkhäuser Verlag, 2001 (Lectures in Mathematics ETH Zürich). – Notes taken by Felix Schlenk. – ISBN 3–7643– 6403–3

[46] Washington, L. C. : Introduction to cyclotomic fields. Second edition. New York: Springer-Verlag, 1997. – Graduate Texts in Mathematics, Bd. 83. – ISBN 0–387–94762–0

[47] Wilson, J. S. : Profinite groups. New York: The Clarendon Press Oxford University Press, 1998 (London Mathematical Society Monographs. New Series. Bd. 19). – ISBN 0–19–850082–3